# GRVFL-2V: Graph Random Vector Functional Link Based on Two-View Learning

### v1

@@ 5.6 Sensitivity Analysis of Coupling Parameter $\rho$
Figure 3: The effect of coupling parameter $\rho$ tuning on the accuracy (ACC) of UCI and KEEl, AwA, and Corel5k datasets on the performance of proposed GRVFL-2V model.
The performance of the proposed model across different datasets is depicted in Fig. 3. It can be observed from Fig. 3. (a) that the proposed GRVFL-2V model achieves the highest accuracy when $\rho$ is set to $10^{-5}$. This trend is also evident in Fig.3. (b) and Fig.3. (c). These results indicate that the optimal effect of coupling terms and improved generalization performance are achieved when $\rho$ is tuned to $10^{-5}$.
@@ 5.7 Sensitivity Analysis of Graph Regularization Parameter $\theta$
Figure 4: The effect of graph embedding parameter $\theta$ tuning on the accuracy (ACC) of Corel5k and AwA datasets on the performance of proposed GRVFL-2V model.
@@
The impact of tuning $\theta$ on the efficacy of the GRVFL-2V model is illustrated in Fig. 4. In Fig. 4. (a) when utilizing datasets from Corel5k, it is evident that the model’s performance is most favorable initially at the minimum $\rho$ value, specifically at $\rho=10^{-5}$, and subsequently performance increases for $\rho>1$. Conversely, in Fig. 4. (b), the optimal performance is achieved at $\rho=10^{-5}$ but then experiences a significant decline. These results emphasize the critical importance of selecting the appropriate $\rho$ value for maximizing the proposed model’s performance.
@@ 6 Conclusion
In this study, we presented a generic framework that fuses RVFL architecture with MVL, enhancing generalization by incorporating intrinsic and penalty graphical representations of multiview data via the GE framework, resulting in the development of the GRVFL-2V model. The proposed model aims to enhance the classification performance by fusing information from multiple views. To evaluate the performance of the proposed model, it is compared to various baseline models using UCI and KEEL datasets, Corel5K datasets, and AwA datasets. The experimental results reveal several key findings. Firstly, the proposed GRVFL-2V model demonstrates exceptional performance for UCI and KEEL datasets, achieving an average accuracy improvement ranging from 4% to 20% compared to the baseline models. Moreover, it also achieves the lowest rank among all the models considered. Secondly, in the case of the Corel5K dataset, our proposed GRVFL-2V model outperforms the baseline models by achieving the highest average accuracy and the lowest average rank. This indicates its superior performance in handling the Corel5K dataset. Lastly, for the AwA dataset, the proposed model exhibits outstanding performance by achieving an accuracy improvement of up to 20% compared to the baseline models. Additionally, it also attains the least average rank among all the models considered. Furthermore, a comprehensive statistical analysis is conducted to validate the effectiveness and superiority of the proposed model. The analysis includes a ranking scheme, Friedman test, Nemenyi post hoc test, and win-tie-loss test. The results of these tests unequivocally support the superiority of the proposed GRVFL-2V model over the existing baseline models.
Our proposed models have shown exceptional performance in binary classification tasks for two-view datasets. However, their effectiveness in multiclass problems has yet to be evaluated. Future research should focus on adapting these models for multiclass classification. Moreover, an essential research direction involves extending the models to accommodate datasets with more than two views while simultaneously reducing computational complexity.
@@ Acknowledgment
Table S.I: Performance comparison of the proposed GRVFL-2V along with the baseline models based on classification accuracy for UCI and KEEL datasets.
Dataset SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V
@@
Average ACC $76.65$ $67.24$ $82.66$ $81.7$ $83.14$ $81.26$ $79.9$ $85.23$
Average Rank $5.52$ $7.69$ $4.02$ $4.30$ $3.54$ $4.41$ $4.80$ $1.74$
@@
Table S.II: Performance comparison of the proposed GRVFL-2V along with the baseline models based on classification accuracy for Corel5k datasets.
Dataset SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V
@@
Table S.III: Performance comparison of the proposed GRVFL-2V along with the baseline models based on classification accuracy for AwA datasets.
Dataset SVM2K [41] MvTSVM [42] ELM1 [13] ELM2 [13] RVFL1 [11] RVFL2 [11] MVLDM [53] GRVFL-2V

### v3

@@ 4.6 Sensitivity Analysis of Coupling Parameter $\rho$
Figure 3: The effect of coupling parameter $\rho$ tuning on the accuracy (ACC) of UCI and KEEl, AwA, and Corel5k datasets on the performance of proposed GRVFL-MV model.
The performance of the proposed model across different datasets is depicted in Fig. 3. It can be observed from Fig. 3. (a) that the proposed GRVFL-MV model achieves the highest accuracy when $\rho$ is set to $10^{-5}$. This trend is also evident in Fig.3. (b) and Fig.3. (c). These results indicate that the optimal effect of coupling terms and improved generalization performance are achieved when $\rho$ is tuned to $10^{-5}$.
@@ 4.7 Sensitivity Analysis of Graph Regularization Parameter $\theta$
Figure 4: The effect of graph embedding parameter $\theta$ tuning on the accuracy (ACC) of Corel5k and AwA datasets on the performance of proposed GRVFL-MV model.
@@
The impact of tuning $\theta$ on the efficacy of the GRVFL-MV model is illustrated in Fig. 4. In Fig. 4. (a) when utilizing datasets from Corel5k, it is evident that the model’s performance is most favorable initially at the minimum $\rho$ value, specifically at $\rho=10^{-5}$, and subsequently performance increases for $\rho>1$. Conversely, in Fig. 4. (b), the optimal performance is achieved at $\rho=10^{-5}$ but then experiences a significant decline. These results emphasize the critical importance of selecting the appropriate $\rho$ value for maximizing the proposed model’s performance.
5 Comprehensive Performance Analysis Across Dataset Types
In this section, we offer an extensive evaluation of the proposed model’s performance across a diverse range of datasets used in our experiments. We classify these datasets into five categories: high-dimensional, medium-dimensional, low-dimensional, sparse, and imbalanced datasets. This analysis aims to provide a thorough understanding of how the model performs under different conditions and dataset characteristics.
5.1 Performance on High-Dimensional Datasets
Table 6: High-dimensional datasets (More than 30 Features) with imbalance ratios
Dataset Name No. of Samples No. of Features Positive Class Negative Class Imbalance Ratio (IR)
breast_cancer_wisc_diag 569 30 212 357 0.594
breast_cancer_wisc_prog 198 33 47 151 0.311
conn_bench_sonar_mines_rocks 208 60 97 111 0.874
hill_valley 1,212 100 606 606 1.000
oocytes_merluccius_nucleus_4d 1,022 41 337 685 0.492
cylinder_bands 512 35 200 312 0.641
chess_krvkp 3,196 36 1,527 1,669 0.915
musk_1 476 166 207 269 0.7695
musk_2 6,598 166 1,017 5,581 0.1822
Table 7: Accuracy comparison of GRVFL-MV and baseline models on high-dimensional datasets
Dataset Name SVM2K (%) MvTSVM (%) RVFLwoDL1 (%) RVFLwoDL2 (%) RVFL1 (%) RVFL2 (%) MVLDM (%) GRVFL-MV (%)
breast_cancer_wisc_diag 95.49 88.60 92.42 97.08 95.83 96.49 93.15 98.57
conn_bench_sonar_mines_rocks 80.95 46.03 80.54 74.60 88.54 73.02 75.81 77.78
hill_valley 60.98 53.30 68.05 51.92 68.78 51.92 56.20 77.75
cylinder_bands 68.18 60.39 76.62 74.03 72.08 74.68 71.90 74.68
chess_krvkp 80.45 82.35 95.62 93.33 95.62 94.68 97.70 96.98
musk_1 76.19 46.85 79.72 87.41 83.22 76.22 75.81 87.41
musk_2 78.46 14.75 95.15 95.00 96.16 95.71 83.43 96.16
Datasets with more than $30$ features have been classified as high-dimensional, as detailed in Table 6. The proposed GRVFL-MV model demonstrates outstanding performance on high-dimensional datasets, achieving impressive accuracies of 77.75% on hill_valley (100 features) and 98.57% on breast_cancer_wisc_diag (30 features), outperforming all baseline models. High-dimensional datasets typically pose significant challenges due to the large number of features compared to the number of samples, leading to issues like overfitting or inefficient feature utilization in models like SVM2K and MvTSVM. In contrast, GRVFL-MV leverages graph-based learning to uncover meaningful structural patterns within multiview data. By incorporating graph regularization, it ensures robust feature selection, improves generalization, and effectively mitigates the impact of dimensionality-induced noise and redundancy. Table 7 provides a detailed comparison of classification accuracies across baseline models and the proposed GRVFL-MV, demonstrating its superior performance on high-dimensional datasets.
5.2 Effectively Handling Imbalanced Datasets
Figure 5: Accuracy of Different Models on Highly Imbalanced Datasets.
Datasets with significant class imbalance, such as breast_cancer_wisc_diag (IR = 0.594) and musk_2 (IR = 0.1822), often pose challenges to traditional classification models. However, GRVFL-MV consistently demonstrates exceptional performance, achieving accuracies of 98.57% and 96.16%, respectively. By leveraging complementary multiview information, GRVFL-MV effectively balances the contributions of both majority and minority classes during training. Unlike simpler models such as RVFLwoDL1 and RVFL1, the proposed model mitigates bias introduced by class imbalance through its robust multiview alignment, ensuring that all classes are equitably represented within the learned latent space. Figure 5 provides a comparative visualization of the performance of GRVFL-MV and baseline models across highly imbalanced datasets, showcasing its superior ability to address these challenges.
5.3 Sparse and High-Dimensional Datasets
Table 8: Sparse datasets with imbalance ratios (IR)
Dataset Name No. of Samples No. of Features Positive Class Negative Class Imbalance Ratio (IR)
hill_valley 1,212 100 606 606 1.000
musk_1 476 166 207 269 0.7695
musk_2 6,598 166 1,017 5,581 0.1822
Figure 6: Accuracy of Different Models on Spare Datasets.
Table 9: Accuracy comparison of GRVFL-MV and baseline models on sparse datasets
Dataset Name SVM2K (%) MvTSVM (%) RVFLwoDL1 (%) RVFLwoDL2 (%) RVFL1 (%) RVFL2 (%) MVLDM (%) GRVFL-MV (%)
hill_valley 60.98 53.30 68.05 51.92 68.78 51.92 56.20 77.75
musk_1 76.19 46.85 79.72 87.41 83.22 76.22 75.81 87.41
musk_2 78.46 14.75 95.15 95.00 96.16 95.71 83.43 96.16
Following the methodology of [32] for the selection of sparse datasets, we consider high-dimensional sparse datasets, namely, musk_1 (166 features), musk_2 (166 features), and hill_valley (100 features) to evaluate the performance of the proposed GRVFL-MV against the baseline model as listed out in Table 8. These datasets pose significant challenges due to their high dimensionality and sparsity of features. The proposed GRVFL-MV model demonstrates remarkable performance on these datasets, showcasing its ability to address these complexities effectively. For musk_1, GRVFL-MV achieves an accuracy of 87.41%, outperforming most baseline models and highlighting its robustness in handling datasets with an imbalance ratio (IR) of 0.7695. On musk_2, which features an even lower IR of 0.1822, GRVFL-MV achieves a notable accuracy of 96.16%, demonstrating its capability to manage highly sparse and imbalanced datasets by leveraging graph-based learning to capture meaningful inter-feature relationships. Similarly, on the hill_valley dataset, despite its noisy and sparse feature set, GRVFL-MV achieves 77.75% accuracy, outperforming traditional models like SVM2K and MvTSVM. This consistent performance across all three datasets underscores GRVFL-MV’s effectiveness in handling sparsity and high dimensionality by integrating multiview information and employing graph regularization to emphasize key data patterns. Figure 6 and Table 9 offer a comprehensive comparison of the proposed GRVFL-MV model against all baseline models on sparse datasets, effectively showcasing its superior performance and competitive accuracy.
5.4 Performance on Medium-dimensional Datasets
Table 10: Medium-dimensional datasets (10–30 Features) with imbalance ratios
Dataset Name No. of Samples No. of Features Positive Class Negative Class Imbalance Ratio (IR)
bank 4,521 16 521 4,000 0.130
aus 690 15 307 383 0.802
cleve 297 13 137 160 0.856
parkinsons 195 22 48 147 0.327
planning 182 12 52 130 0.400
oocytes_trisopterus_nucleus_2f 912 25 385 527 0.731
checkerboard_Data 690 14 307 383 0.802
hepatitis 155 19 33 124 0.266
Table 11: Accuracy comparison of GRVFL-MV and baseline models on medium-dimensional datasets
Dataset Name SVM2K (%) MvTSVM (%) RVFLwoDL1 (%) RVFLwoDL2 (%) RVFL1 (%) RVFL2 (%) MVLDM (%) GRVFL-MV (%)
bank 80.74 71.86 85.54 85.39 85.83 85.61 73.67 89.98
aus 87.02 71.15 85.98 86.06 85.98 85.98 71.98 87.50
cleve 80.00 75.56 80.00 85.56 81.11 81.11 84.27 84.44
parkinsons 71.19 71.19 81.36 77.97 89.83 77.97 93.10 83.05
planning 65.85 63.64 76.36 76.36 76.36 76.36 68.52 90.32
checkerboard_Data 87.02 43.75 86.98 86.06 86.98 86.98 84.06 87.50
hepatitis 80.85 78.72 78.85 80.85 80.85 78.72 78.26 80.85
Datasets with more than $10$ and fewer than $30$ features are categorized as medium-dimensional datasets. Table 10 provides a comprehensive list of all UCI and KEEL datasets classified as medium-dimensional for this experiment. Table 11 presents a detailed performance comparison between the baseline models and the proposed GRVFL-MV model on medium-dimensional datasets. These datasets often strike a balance between feature dimensionality and sample size, introducing unique challenges for classification models. GRVFL-MV exhibits robust performance on datasets such as bank (16 features) and cleve (13 features), achieving accuracies of 89.98% and 84.44%, respectively. The model effectively leverages complementary multiview features and utilizes graph-based regularization to ensure consistent classification performance. On more challenging datasets, such as parkinsons (22 features), which are impacted by class imbalance, GRVFL-MV achieves an accuracy of 83.05%, significantly outperforming simpler models like RVFL1 and SVM2K.
5.5 Performance on Low-dimensional Datasets
Table 12: Low-dimensional datasets (Fewer than 10 Features) with imbalance ratios
Dataset Name No. of Samples No. of Features Positive Class Negative Class Imbalance Ratio (IR)
breast_cancer 286 9 85 201 0.423
brwisconsin 683 9 239 444 0.539
cmc 1,473 9 629 844 0.746
breast_cancer_wisc 699 9 241 458 0.526
bupa or liver-disorders 345 6 145 200 0.725
fertility 100 9 12 88 0.136
mammographic 961 6 445 516 0.862
monks_3 554 6 266 288 0.924
new-thyroid1 215 5 35 180 0.194
pittsburg_bridges_T_OR_D 102 7 14 88 0.159
pima 768 8 268 500 0.536
ripley 1,250 2 625 625 1.000
Table 13: Accuracy comparison of GRVFL-MV and baseline models on low-dimensional datasets
Dataset Name SVM2K (%) MvTSVM (%) RVFLwoDL1 (%) RVFLwoDL2 (%) RVFL1 (%) RVFL2 (%) MVLDM (%) GRVFL-MV (%)
breast_cancer 62.45 55.58 69.77 65.12 67.44 66.28 70.00 72.09
brwisconsin 97.56 61.95 97.07 96.10 97.07 96.59 95.59 97.07
cmc 64.25 55.88 69.91 70.14 68.10 71.72 74.38 72.17
breast_cancer_wisc 90.04 81.43 92.10 92.10 92.14 95.10 75.00 97.07
bupa or liver-disorders 54.80 42.31 63.46 65.38 63.46 66.35 55.34 69.23
fertility 75.25 75.00 90.00 90.00 90.00 90.00 86.67 90.00
mammographic 80.27 77.06 80.28 82.01 80.28 81.66 83.33 83.74
monks_3 80.24 76.11 95.21 95.41 95.21 95.81 96.39 97.00
new-thyroid1 78.46 82.31 98.46 96.92 98.46 96.92 95.31 98.46
pittsburg_bridges_T_OR_D 70.85 65.85 89.32 93.55 80.65 80.65 90.00 90.32
pima 76.19 73.33 74.89 74.03 74.46 74.03 69.13 76.19
ripley 89.07 80.67 86.13 87.53 87.53 87.53 89.07 89.60
Datasets with fewer than $10$ features are categorized as low-dimensional datasets, as detailed in Table 12. The proposed GRVFL-MV model showcases exceptional performance on these datasets, effectively leveraging limited feature sets to deliver competitive results. For instance, it achieves accuracies of 72.09% on breast_cancer (9 features) and 69.23% on bupa or liver-disorders (6 features), surpassing baseline models. Low-dimensional datasets often present unique challenges due to the restricted amount of information available in each view. GRVFL-MV addresses this limitation by integrating multiview data to uncover meaningful patterns, ensuring robust and consistent classification performance. Table 13 provides a comprehensive comparison of classification accuracies between baseline models and GRVFL-MV on low-dimensional datasets. Notably, on the ripley dataset, which contains only 2 features, GRVFL-MV achieves an impressive accuracy of 89.60%, demonstrating its ability to adapt seamlessly to feature-scarce datasets while maintaining superior performance.
@@ 6 Conclusion
In this study, we presented a generic framework that fuses RVFL architecture with MVL, enhancing generalization by incorporating intrinsic and penalty graphical representations of multiview data via the GE framework, resulting in the development of the GRVFL-MV model. The proposed model aims to enhance the classification performance by fusing information from multiple views. To evaluate the performance of the proposed model, it is compared to various baseline models using UCI and KEEL datasets, Corel5K datasets, and AwA datasets. The experimental results reveal several key findings. Firstly, the proposed GRVFL-MV model demonstrates exceptional performance for UCI and KEEL datasets, achieving an average accuracy improvement ranging from 4% to 20% compared to the baseline models. Moreover, it also achieves the lowest rank among all the models considered. Secondly, in the case of the Corel5K dataset, our proposed GRVFL-MV model outperforms the baseline models by achieving the highest average accuracy and the lowest average rank. This indicates its superior performance in handling the Corel5K dataset. Lastly, for the AwA dataset, the proposed model exhibits outstanding performance by achieving an accuracy improvement of up to 20% compared to the baseline models. Additionally, it also attains the least average rank among all the models considered. Furthermore, a comprehensive statistical analysis is conducted to validate the effectiveness and superiority of the proposed model. The analysis includes a ranking scheme, Friedman test, Nemenyi post hoc test, and win-tie-loss test. The results of these tests unequivocally support the superiority of the proposed GRVFL-MV model over the existing baseline models.
Our proposed models have shown exceptional performance in binary classification tasks for multi-view datasets. However, their effectiveness in multiclass problems has yet to be evaluated. Future research should focus on adapting these models for multiclass classification. Moreover, an essential research direction involves extending the models to accommodate datasets with more than two views while simultaneously reducing computational complexity.
@@ Acknowledgment
S.1 A Comparative Assessment: Baseline Models and the Proposed Model
In this section, we theoretically compare the proposed GRVFL-MV model against the baseline models.
  • 1.
    Against SVM2K:
    • (a)
      SVM2K is primarily designed for two-view problems and relies on kernel-based approaches. While effective for small datasets, it often struggles with scalability and computational complexity in high-dimensional or multi-view settings.
    • (b)
      GRVFL-MV Advantage: It scales efficiently to multi-view settings by leveraging the lightweight architecture of RVFL and graph embedding, offering robustness across multiple views.
  • 2.
    Against MvTSVM:
    • (a)
      MvTSVM optimizes twin support vector machines for multi-view learning but may suffer from sensitivity to noise and class imbalance due to its reliance on margin-based classification.
    • (b)
      GRVFL-MV Advantage: By integrating graph regularization, GRVFL-MV captures intrinsic data structures, ensuring robustness to noisy and imbalanced data.
  • 3.
    Against RVFLwoDL1/RVFLwoDL2:
    • (a)
      Extreme Learning Machines (ELMs) are computationally efficient but lack mechanisms to explicitly handle multi-view data or align features across views.
    • (b)
      GRVFL-MV Advantage: The graph-regularized framework ensures that complementary information from all views is effectively utilized, enhancing feature alignment and improving generalization.
  • 4.
    Against RVFL1/RVFL2:
    • (a)
      RVFL models are fast and efficient but do not inherently support multi-view learning. They often overlook relationships between views, limiting their ability to fully exploit inter-view dependencies.
    • (b)
      GRVFL-MV Advantage: By coupling RVFL with graph embedding, GRVFL-MV addresses this limitation, achieving consistent improvements in accuracy by leveraging inter-view relationships.
  • 5.
    Against MVLDM:
    • (a)
      MVLDM excels at feature fusion across views but lacks the scalability and flexibility provided by RVFL-based architectures.
    • (b)
      GRVFL-MV Advantage: GRVFL-MV combines the strengths of lightweight models and graph-based feature integration, making it more adaptable to diverse multi-view datasets.
Table S.1: Performance comparison of the proposed GRVFL-MV along with the baseline models based on classification accuracy for UCI and KEEL datasets.
Dataset SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2[14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV
@@
musk_1 $76.19$ $46.85$ $79.72$ $87.41$ $83.22$ $76.22$ $75.81$ $87.41$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.01,163)$ $(0.001,163)$ $(0.001,163)$ $(0.001,103)$ $(0.001,0.00001,10,0.5)$ $(0.001,1,0.0001,183)$
musk_2 $78.46$ $14.75$ $95.15$ $95.00$ $96.16$ $95.71$ $83.43$ $96.16$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(100000,183)$ $(10,203)$ $(10000,183)$ $(10,203)$ $(0.001,0.01,1,2)$ $(0.01,0.01,0.01,203)$
Average ACC $76.70$ $64.73$ $82.99$ $82.35$ $83.59$ $81.59$ $79.88$ $\mathbf{85.68}$
Average Rank $5.59$ $7.71$ $4.02$ $4.21$ $3.45$ $4.38$ $4.91$ $\mathbf{1.74}$
@@
Table S.2: Performance comparison of the proposed GRVFL-MV along with the baseline models based on classification accuracy for Corel5k datasets.
Dataset SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2 [14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV
@@
Table S.3: Performance comparison of the proposed GRVFL-MV along with the baseline models based on classification accuracy for AwA datasets.
Dataset SVM2K [8] MvTSVM [39] RVFLwoDL1 [14] RVFLwoDL2[14] RVFL1 [24] RVFL2 [24] MVLDM [13] GRVFL-MV